The amount of time in minutes that it took to complete a grocery store checkout for a sample of 5 transactions is shown below:
Checkout time- 2, 11, 1, 6, 5



What is the variance?
Answer #1

mean = (2 + 11 + 1 + 6 + 5) / 5 = 5

Sample variance = [(2 - 5)2 + (11 - 5)2 + (1 - 5)2 + (6 - 5)2 + (5 - 5)2] / 4 = 15.5 (ans)